Y=ƒ(2x-1)/(x^2+1)and ƒ'(x)= sinx^2 then dy/dx


Note x^2 means  x square

Grade:

1 Answers

student iit
14 years ago

I guess you mean dY/dx?

If f'(x)= sinx^2 then dy/dx= sinx^2 => dy=sinx^2 dx , integrate to get the function y=f(x).Now calculate f(2x-1) and then do dY/dx.